Please Help Me Anser Following A-D :)

A)Find the area of the shaded region. The graph to the right depicts IQ scores of adults, and those scores are normally distrib

B)Find the area of the shaded region. The graph to the right depicts IQ scores of adults, and those scores are normally distribC)Find the area of the shaded region. The graph to the right depicts IQ scores of adults, and those scores are normally distribD)Find the indicated IQ score. The graph to the right depicts IQ scores of adults, and those scores are normally distributed wi
Answer #1

H=100, 5=159 Z. KU, X-100 @ PlQ3125)= Plz=125-400) =P(Z < 25)- Plz41.67) 18.9525 LX4-125 P- 24 © Play92) = PLZ > 92-100) - Pl